Find the surface area of a right prism with equalaterial trianglar bases if all edges are two units long.
Answer by sarah_adam(201)

Triangular based prism with equilateral triangle as base:
For a prism with equilateral triangle as base
h1 is the height of the triangle
at the base ( or the top )
h is height of the prism
h1 = (3^(1/2) / 2 )* a = sqrt(3)/2 *a where a is the length of side
h1 = sqrt(3)/2 * 2 = sqrt(3)
Surface area of Prism: a.h1 + h
SA = 2(sqrt(3))+ h